# Designs for Enabling Collaboration in Human-Machine Teaming via Interactive and Explainable Systems

### v1

@@ 4.2 Modifying an Interpretable Policy
While the above architecture can be used alongside RL to produce a collaborative AI policy, the result may not actually be helpful or what the human wants. Humans, when teaming with machines, should be able to intuitively update what the robot has learned or change it based upon preferences that may evolve over time. Such is critical in the positive development of coordination strategies and is associated with the calibration of trust, assignment of roles, and development of a shared mental model. As such, we propose a policy modification scheme that allows the user to repeatedly team with an AI maintaining an IDCT policy, visualize the current behavior in tree form, and modify its AI’s behavior. The iterative process generated through this scheme can facilitate a feedback loop, allowing for the possibility of team development and improved HMT performance over teaming episodes.
We term our modification scheme human-led policy modification. We provide humans with an explicit pathway to “communicate” with an AI after each teaming interaction through a GUI, with capabilities displayed in Figure 3. Within this interface, users start with the pre-trained collaborative AI IDCT policy and can modify the AI’s behavior by creating a new tree structure that may vary in what state features appear in the decision nodes, actions taken in leaf nodes, and the respective probabilities of actions within the leaf node. It is important to note that users are limited to expanding the tree to a depth of four (i.e., a max of 16 leaves).
Figure 3: Users have several capabilities in creating an effective teammate, including modifying the tree structure by adding or removing decision nodes (top), changing state features the tree is conditioned on (left), and modifying actions and/or their respective probabilities at leaf nodes (right).
@@ 5 Human-Subjects Study
Here, we discuss our between-subjects user study that seeks to understand how users interact with an AI across repeated play under different factors. Below, we introduce our research questions, provide a description of the independent variables and procedure, include brief descriptions of the behaviors learned by collaborative AI, and finally discuss our findings.
@@
    RQ1: How does HMT performance vary across factors?
@@
    IV1-C2: AI-Led Policy Modification: After interacting with the agent, the AI utilizes recent gameplay to fine-tune a human gameplay model via Behavioral Cloning and performs reinforcement learning for five minutes to optimize its own policy to better support the human teammate. Upon completion of policy optimization, the user can visualize the updated AI policy in its interpretable tree form prior to the next interaction. This is similar to HA-PPO Carroll et al. (2019), adapted to an online setting.We limit the online optimization time for the AI teammate to five minutes to create a feasible user-study. Online optimization via RL becomes challenging as a limited samples can be obtained in this time, and the policy is not guaranteed to improve. In cases where the policy degrades, we use the original policy prior to optimization.
@@
    IV1-C5: Static Policy - Fictitious Co-Play: Strouse et al. (2021): User teams with an AI maintaining a static black-box, neural network (NN) policy trained across a diverse partner set. As this is a baseline, we utilize an NN rather than the human-understandable IDCT policy model used in all other conditions (IV1:C1-4).
@@
For IV2, we consider the following domains in Figure 4:
@@
    IV2-D1: Forced Coordination: Users team with an AI that is separated by a barrier and must pass over onions and plates in a timely manner. In this domain, agents are forced to collaborate.

### v2

@@ 4.2 Modifying an Interpretable Policy
While the above architecture can be used alongside RL to produce a collaborative AI policy, the result may not actually be helpful or what the human wants. Humans, when teaming with machines, should be able to intuitively update what the robot has learned or change it based upon preferences that may evolve over time. Such is critical in the positive development of coordination strategies and is associated with the calibration of trust, assignment of roles, and development of a shared mental model. As such, we propose a policy modification scheme that allows the user to repeatedly team with an AI maintaining an IDCT policy, visualize the current behavior in tree form, and modify its AI’s behavior.
Figure 3: Users have several capabilities in creating an effective teammate, including modifying the tree structure by adding or removing decision nodes, changing state features the tree is conditioned on, and modifying actions and/or their respective probabilities at leaf nodes.
The iterative process generated through this scheme can facilitate a feedback loop, allowing for the possibility of team development and improved HMT performance over teaming episodes.
We term our modification scheme human-led policy modification. We provide humans with an explicit pathway to “communicate" with an AI after each teaming interaction through a GUI, with capabilities displayed in Figure 3. Within this interface, users start with the pre-trained collaborative AI IDCT policy and can modify the AI’s behavior by creating a new tree structure that may vary in what state features appear in the decision nodes, actions taken in leaf nodes, and the respective probabilities of actions within the leaf node. It is important to note that users are limited to expanding the tree to a depth of four (i.e., a max of 16 leaves), and the modification is not timed.
4.3 Trained Collaborative Teammate Policies
Across our experiment, we study collaboration in two domains, Forced Coordination and Optional Collaboration, displayed on the left-hand side of Figure 4. In each domain, we train an IDCT policy via agent-agent collaborative training and a neural network (NN) policy following the population-based training scheme in Strouse et al. [40]. In the first domain of Forced Coordination, the IDCT policy converged to a policy with an average reward of $315.22\pm 14.59$, and the neural network policy converged to an average reward of $403.16\pm 16.08$ evaluated over 50 teaming simulations with the synthetic human teammate the policy was trained with. In the second domain, Optional Collaboration, the IDCT policy converged to a policy with an average reward of $171.46\pm 18.89$, and the neural network policy converged to an average reward of $295.02\pm 1.86$. Thus, a consequent confound due to the current difference in performance capabilities between interpretable vs. black-box models is that the NN policy outperforms the IDCT policy in both domains. This displays a need for improving optimization algorithms for interpretable models representing collaborative agent policies. However importantly, while the initial simulated performance of interpretable models may underperform black-box models, the ability for humans to understand machine behavior and improve upon behavior may allow these approaches to compete or even outperform black-box NN models. We can also compare to the heuristic policies presented in Section 3, observing that the training performance of the IDCT and NN policies in the Optional Collaboration domain underperform the collaborative heuristic (408 vs. 295.02 and 171.46). We provide visualizations of the trained IDCT policies for each domain in the appendix, finding that after contextual pruning, the AI IDCT policy has two and three leaves, respectively.
@@ 5 Human-Subjects Study
Here, we discuss our between-subjects user study that seeks to understand how users interact with an AI across repeated play under different factors. Below, we introduce our research questions, provide a description of the independent variables and procedure, and discuss our findings.
@@
    RQ1: How does human-machine teaming performance vary across factors?
@@
    IV1-C2: AI-Led Policy Modification: After interacting with the agent, the AI utilizes recent gameplay to fine-tune a human gameplay model via Behavioral Cloning and performs reinforcement learning for five minutes [^2] We limit the online optimization time for the AI teammate to five minutes to create a feasible user-study. This RL optimization is challenging as only a limited number of samples can be obtained in this time, and thus, the policy is not guaranteed to improve. In cases where the policy degrades, we use the original policy prior to optimization. to optimize its own policy to better support the human teammate. Upon completion of policy optimization, the user can visualize the updated AI policy in its interpretable tree form prior to the next interaction. This is similar to HA-PPO [5], adapted to an online setting.
@@
    IV1-C5: Static Policy - Fictitious Co-Play: [40]: User teams with an AI maintaining a static black-box, neural network (NN) policy trained across a diverse partner set. As this is a baseline, we utilize an NN rather than the legible IDCT policy used in other conditions (IV1:C1-4).
For IV2, we consider the following domains displayed on the left-hand side of Figure 4:
  1. 1.
    IV2-D1: Forced Coordination: Users team with an AI that is separated by a barrier and must pass over items in a timely manner. Here, agents are forced to collaborate.
  2. 2.
@@
Procedure: A participant is first randomly placed into one of the five conditions in IV1. The participant starts with a pre-experiment survey collecting demographic information, experience with video games and decision trees, and the Big Five Personality Questionnaire [7]. Afterward, a participant conducts a brief tutorial in Overcooked with a random AI agent, improving the user’s understanding of game controls and the assigned task. Once completed, the primary experimentation begins. Users will team with an AI four times in each domain (randomly ordered), starting with the unique domain-specific pre-trained agent, and are told that their goal is to maximize their score in the last teaming interaction, the “performance round." After each teaming interaction, in the first three factors, the user will modify and visualize the AI’s policy (IV1-C1), the AI will optimize its own policy proceeded by user visualization (IV1-C2), or the user will solely view the policy (IV1-C3). In IV1-C4 and IV1-C5, as the AI is black-box (perceived to be black-box in IV1-C4 and truly black-box in IV1-C5), transitionary pages are shown to the participant, providing them a pause before they team with the agent again. Upon completion of the condition-specific (or lack of) actions, users complete a NASA-TLX Workload Survey. After users have completed a domain, providing us with four episodes of teaming data and workload assessments, we administer several post-study scales, including the Human-Robot Collaborative Fluency Assessment [15], Inclusion of Other in the Self scale [1], and Godspeed Questionnaire [2]. Upon completion of the two domains, the experiment concludes.
(a) Performance Data in IV2-D1: Forced Coordination.
(b) Performance Data in IV2-D2: Optional Collaboration
Figure 4: User gameplay scores across teaming iterations with per-iteration means connected by the red dotted line and the per-iteration standard deviation shaded in red.
5.1 Results
Our experiment is a 5 (teaming method; between-subjects) $\times$ 2 (no. of domains; within-subjects) $\times$ 4 (no. of repeated evaluations; within-subjects) mixed-factorial experiment. We recruited 50 participants under an IRB-approved protocol, whose ages range from 18 to 32 (Mean age: 24.14; Std. Dev.: 4.10; 46% Female, 52% Male, 2% Non-Binary), with participants randomly assigned to each of the factor levels, with ten total subjects per level. The duration of the experiment was $70.98\pm 19.71$ minutes  [^3] The significant variance in experiment duration arises from the granularity across our conditions. The increase in human effort to understand and interact with the policy results in an increase in duration. We note that as our experiment is relatively short, it is unlikely that experiment fatigue played a role in our results as would be common in experiments with large task variances.. Our data was modeled as a full-factorial, between-subjects ANOVA. We test for normality and homoschedascity (see appendix) and employed a corresponding non-parametric test if the data failed to meet these assumptions. We display our objective findings in the right-hand side of Figure 4.
RQ1: Team Performance: In analyzing reward, we find trends with respect to the maximum reward participants obtained within a domain across iterations (Figure 5). Using Friedman’s test, we find a significant difference across domains ($\chi^{2}$(1)=46.08, $p<0.001$) and analyze the domains separately.
In IV2-D2, a Kruskal-Wallis Test was conducted to analyze differences in participant teaming performance across conditions, finding a significant effect ($\chi^{2}(4)=29.922,p<0.001$). We conduct post-hoc pairwise comparisons, utilizing Dunn’s test, and find that IV1-C5 (Ficticious Co-Play) is significantly better than IV1-C2 ($p<0.001$), IV1-C3 ($p<0.001$), and IV1-C4 ($p<0.001$), and IV1-C1 (Human-Led Policy Modication) is significantly better than IV1-C2 ($p<0.05$), IV1-C3 ($p<0.05$), and IV1-C4 ($p<0.05$). For white-box AI teammates (IV1:C1-3), the latter finding displays the benefit of Human-Led Policy Modification in improving HMT performance for interpretable models. These findings display that 1) white-box approaches supported with policy modification can outperform white-box approaches alone, 2) black-box models can outperform white-box approaches in HMT, and 3) by comparing IV1-C3 to IV1-C4, interpretability alone afforded via tree visualizations did not provide any direct objective benefits. Finally, in Optional Collaboration, across all conditions we see that HMT scores are not near that of the collaborative heuristic, displaying a gap that must be addressed to achieve effective HMT.
RQ2: Team Development: In analyzing RQ2, we look at the change in reward across iterations one to four and relate our findings to Tuckman’s model. Utilizing a Friedman’s test, we find a difference across domains ($\chi^{2}$(1)=20.48, p$<$0.001) and analyze the domains separately. In IV2-D1, we see that none of the conditions results in a significant improvement in teaming performance over repeated iterations. In IV2-D2, we see IV1-C1 ($p<0.01$) and IV1-C2 ($p<0.01$) significantly improve over repeated teaming interactions. The improving interactions can be connected to the Norming stage in team development, where teams begin to develop a strategy and team mental models. We see conditions that facilitate Norming have the attribute of policy adaptation and are white-box.
Next, we analyze whether different person-specific factors allow HMT to improve more quickly than
Figure 5: Maximum Reward and Subjective Ratings Across IV1 Factors.
others. In IV2-D1, we find that conscientiousness is trending in its correlation with improvement ($0.10.05<p<$ 0.1). In IV2-D2, we find that participants with high familiarity with Trees improve more across iterations ($F(1)=7.448,p<0.01$). These findings signify that positive interaction with interpretable models may be more beneficial to those with an engineering background and specific personality traits.
Finally, we detect an interesting trend in IV2-D1 under the IV1-C1 condition. We see a drop in performance between the first teaming iteration and later iterations, followed by a rise. We believe this relates to the Forming and Storming stages, where team members are still developing effective strategies to coordinate. As the last iteration shows an improvement in performance, we hypothesize that the team was shifting into the Norming stage. In future, it would be interesting to evaluate a larger number of iterations to see if the behavior would continue to uptrend. This requires further research due to the additional resources and time needed for more teaming iterations.
Subjective Findings: In IV2-D1, we find that users did not find any subjective differences toward the teaming interaction across conditions. In IV2-D2 (Figure 5), we find that users find collaboration with AIs under condition IV1-C2 and IV1-C4, on average as less fluent than IV1-C1 (p$<$0.01, p$<$0.01), and IV1-C4 as less fluent than IV1-C5 ($p<0.05$). Users also trusted the AI and perceived the AI contributed more in IV1-C5 than in IV1-C2 (p$<$0.05, p$<$0.05) and IV1-C4 ($p<0.05,p<0.05$). Furthermore, the users viewed the AI more positively in IV1-C1 and IV1-C5 than in both IV1-C2 (p$<$0.05, p$<$0.05) and IV1-C4 (p$<$0.05, p$<$0.01). Overall, participants generally assessed higher-performing agents more positively in their subjective ratings. In considering conditions that utilized a tree-based model (IV1-C1, IV1-C2, IV1-C3, and IV1-C4), we see the addition of interaction with the tree policy provides significant subjective benefits in positive teaming traits and collaborative fluency (defined within the Human-Robot Collaborative Fluency Assessment [2]). In including the remaining condition, which utilizes a black-box model, IV1-C5: Fictitious Co-Play, and comparing it to IV1-C1: Human-Led Policy Modification, we see that even though Fictitious Co-Play outperformed Human-Led Policy Modification in terms of team reward (though not significantly in the domain of Optional Collaboration), no significant subjective differences were observed between these two conditions. This presents an interesting relationship between transparency, interaction, and performance in relation to subjective perception that warrants future research.
Design Guidelines: To achieve fluent HMT, we specify the following forward-facing guidelines.
@@ 5.1 Results
    The creation of white-box learning approaches that can produce interpretable collaborative agents that achieve competitive initial performance to that of black-box agents. This guideline is critical to providing humans with the subjective benefits obtained from interactivity with white-box models, objective benefits of black-box models, and the ability to interact with policies to facilitate team development.